Vanguard Corporation has provided the following data from its activity-based costing system:

Activity Cost Pool Total Cost Total Activity
Assembly $ 999,120 69,000 machine-hours
Processing orders 81,175 1,700 orders
Inspection 122,311 1,730 inspection-hours

The company makes 490 units of product O37W a year, requiring a total of 740 machine-hours, 30 orders, and 10 inspection-hours per year. The product's direct materials cost is $36.52 per unit and its direct labor cost is $29.86 per unit.

According to the activity-based costing system, the average cost of product O37W is closest to:

Multiple Choice

  • $66.38 per unit.

  • $92.61 per unit.

  • $91.17 per unit.

  • $89.69 per unit.
Answer #1

Answer:

Cost per unit Total
Units produced 490
Machine hours required 740 $14.48 =(999120/69000) 10715.2
Orders required 30 $47.75 =(81175/1700) 1432.5
Inseption hours 10 $70.7 =(122311/1730) 707
Materials $36.52 17894.8
Labor $29.86 14631.4
45380.9

Total cost of 490 units of 45380.9
cost per unit = $45380.9/490 = $92.61 per unit